Dimensions of the board? 4 on bottom Strains and stresses (bending stress) were recorded &/or calculated as follows: Gage Strain( Stress (psi Strain (ue) Stress (psi 81 -23 318 -323 795 795 3120 3120 5440 5440 4 154 16. (12 points): Use the above data to complete the following: a. (6 points) Determine an estimate of the Elastic Modulus for this material b. (6 points) Calculate an estimate of Poissons ratio for this material
Answer #1

Concept -We know in a elastic region

STRESS=ELASTICMODULAS\times STRAIN ............................1

  PoissonRatio=- \frac{lateral strain}{axial strain}...............................................................2

So here in a diagram

1,3,5 gauges measure axial stress and strain

4 gauge measure axial compressive stress and strain

2,6 gauges measure lateral stress and strain

Solution

a)using equation 1 we get

Stress= Elasticmodulas* Strain

795=Elasticmodulas*81

Elasticmodulas=795/81=9.81 psi

b)poisson ratio= - strain at 1/strain at 2

=-(-23/81)=0.283 ANS